# Caching-Augmented Lifelong Multi-Agent Path Finding

### v1

@@ Abstract
Multi-Agent Path Finding (MAPF), which involves finding collision-free paths for multiple robots, is crucial in various applications. Lifelong MAPF, where targets are reassigned to agents as soon as they complete their initial objectives, offers a more accurate approximation of real-world warehouse planning. In this paper, we present a novel mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new map grid type called cache for temporary item storage and replacement and designed a lock mechanism for it to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. We identified three main factors significantly impacting CAL-MAPF performance through experimentation: suitable input task distribution, high cache hit rate, and smooth traffic. Overall, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions, maps and agent configurations.
@@ I INTRODUCTION
Automated warehouses represent a multi-billion-dollar industry led by corporations such as Amazon, Alibaba, Ocado, and inVia. These facilities deploy hundreds of robots to transport goods from one location to another [1]. Planning collision-free solutions for hundreds of robots is a key component of these automated warehouses, and it can be simplified to a Multi-Agent Path Finding (MAPF) problem [2]. The MAPF problem requires planning collision-free paths for multiple agents from their start locations to pre-assigned target locations in a known environment while minimizing a given cost function. Many algorithms have been developed to solve this problem optimally and bounded suboptimally, such as Conflict-Based Search (CBS) [3], $M^{*}$ [4] and Enhanced CBS (ECBS) [5].
@@
While many studies focus on developing effective algorithms, there is also interest in enhancing the throughput of automated warehouses within the Lifelong MAPF framework by optimizing warehouse layouts. Prominent approaches include Fishbone design [12], DSAGE [13], and GGO [14]. However, these works mainly focus on static storage design and do not consider input distributions with specific patterns that may shift over time. To address this issue, some studies have explored intelligent decision-making strategies [15, 16, 17]. Notable examples include the Kiva System and various policies like the velocity-based policy [16], and mixed-integer optimization and heuristic methods [18]. Yet, these methods are designed for specific warehouse systems, making it difficult to apply them to different warehouse types. They also require implementing complex rearrangement policies for the entire warehouse storage.
In this paper, we present a dynamic cache delivery mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), inspired by cache design, a foundational element extensively utilized in computer architecture, databases, and various other domains. This cache mechanism is adaptable to diverse Lifelong MAPF algorithms, warehouse storage strategies, and incoming task distributions. We have developed a novel map grid type that allows items to be temporarily stored and replaced. Additionally, we design a lock mechanism to enhance the stability of the planning solution. We have evaluated this cache mechanism using different cache replacement policies and a range of input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: input task distribution, cache hit rate, and map design. Overall, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.
@@ II Problem Definition
Lifelong Multi-Agent Path Finding (Lifelong MAPF) problem presents unique challenges due to the dynamic and unending nature of the tasks. Following the terminology in the MAPF community, we use agents to refer to robots. Let $I=\{1,2,\cdots,N\}$ denote a set of $N$ agents. $G=(V,E)$ represents an undirected graph, where each vertex $v\in V$ represents a possible location of an agent in the workspace, and each edge $e\in E$ is a unit-length edge between two vertices that moves an agent from one vertex to the other. Self-loop edges are allowed, which represent “wait-in-place” actions. Each agent $i\in I$ has a unique start location $s_{i}\in V$. Assume there is an task queue $Q=[q_{1},q_{2},...]$. For each task $q_{i}\in Q$, $q_{i}$ represents a certain item located in a reachable vertex in $V$ that at least one agent can find a path to the vertex. Each item must be delivered to a vertex in $V$ as user specifies. These tasks are generated from a specific probability distribution to enhance realism.
@@
We study an online setting in which incoming tasks are not known in advance. It is assumed that there is a task assigner (external to low-level MAPF algorithm) from which agents can request target locations throughout the system’s operation. The task assigner determines the specific target locations for agents based on tasks in $Q$. In our paper, this task assigner is considered naive and will return target location based on the first unfinished task in $Q$. Each agent $i\in I$ starts from a start location $s_{i}\in V$ and its target location depends on the task assigner. As agents complete their paths at their assigned goals, they are immediately assigned new targets, reflecting warehouse operations’ perpetual and iterative nature. Each action of agents, either waiting in place or moving to an adjacent vertex, takes a time unit. An agent’s path, $p^{i}=[v_{0}^{i},v_{1}^{i},\ldots,v_{T_{i}}^{i}]$, tracks the sequence of vertices traversed by agent $i$ from its start to its target.
As shown in  Figure 1, we primarily focus on 2D warehouse layout maps and categorize all non-obstacle map grids into three types: blue grids $S=\{s_{i}|s_{i}\in V\}$ for shelves that store items, white grids $W=\{w_{i}|w_{i}\in V\}$ for normal aisle grids, and green grids $U=\{u_{i}|u_{i}\in V\}$ for unloading ports where agents deliver items to complete a task. Agents can enter $S$ positions only from $W$ positions. Direct transitions between any two positions in $S$ are not allowed.
@@
We will also involve a new type of map grids: caches. These additional vertices, designated as $C=\{c_{i}|c_{i}\in V\}$ with purple color, serve as interim storage areas to streamline the fulfillment of tasks by reducing the travel time for agents retrieving items. To the best of our knowledge, no previous MAPF works have explored this new cache mechanism, so we adopt the following simplifying assumptions: (1) The warehouse is considered to have an unlimited supply of items, eliminating worries about stock running out. (2) Agents have infinite capacity for carrying items. However, they are limited to transporting one type of item at a time, enabling them to move any quantity of that specific item without restriction. (3) Items removed or evicted from the cache are treated as if they vanish immediately.
@@ III-A MAPF
(One-Shot) MAPF, which has been proved an NP-hard problem with optimality [6], has a long history [19]. It has inspired a wide range of solutions for its related challenges. Decoupled strategies, as outlined in [19, 20, 21], approach the problem by independently planning paths for each agent before integrating these paths. In contrast, coupled approaches [22, 23] devise a unified plan for all agents simultaneously. There also exist dynamically-coupled methods [3, 24] that consider agents that are planned independently at first and then together only when needed in order to resolve agent-agent conflicts. Among these, Conflict-Based Search (CBS) algorithm [3] stands out as a centralized, optimal solution for MAPF, with several bounded-suboptimal variants such as ECBS [5] and EECBS [25]. Some suboptimal MAPF solvers, such as Prioritized Planning (PP) [26, 19], PBS [27], PIBT [10] and their variant methods [28, 29, 11, 30] exhibit better scalability and efficiency.
@@ III-B Lifelong MAPF
Compared to the MAPF problem, Lifelong MAPF continuously assigns new target locations to agents once they have reached their current targets. It is not necessary for agents to arrive at their targets simultaneously in Lifelong MAPF. There are mainly three kinds of methods: solving the problem as a whole [31], using MAPF methods but replanning all paths at each specified length timestep [8, 10], and replanning only when agents reach their current target locations and are assigned new targets [32, 33, 11, 30]. Among them, MAPF-LNS [9, 29] and LaCAM [11, 30] are the leading methods with high task delivery throughput.
@@ III-C Cache
Cache [34] serves as a crucial hardware component designed to temporarily store data, enhancing the efficiency of future request processing. Its main goal is to speed up access to frequently used data, thus minimizing dependence on slower storage disks. Popular caching policies include Least Recently Used (LRU) [35] and First-In-First-Out (FIFO) [36], with studies indicating LRU’s superiority over FIFO [37, 38, 39]. The caching concept is widely applied in various fields, such as Databases [40] and Content Delivery Networks (CDN) [41].
@@ III-D Warehouse Storage Strategy
Automated Storage and Retrieval Systems (AS/RS) have attracted attention for their potential to enhance warehouse efficiency and reduce operational costs [42, 43]. Various strategies for assigning items to storage locations have been widely adopted and evaluated [43, 44]. The random storage policy allocates each item type to a randomly chosen storage location and offers high space utilization [45]. The closest open location storage policy places the arriving unit of items in the nearest available storage location, minimizing the immediate travel distance [46, 47]. The turnover-based storage policy assigns storage locations for each item based on their demand frequency [16, 17].
@@ IV-B Algorithm
As shown in Algorithms 1, 2 and 3, the whole procedure is mainly based on the task assigner which encounters new tasks and decides what locations need to be visited by agents. As illustrated in Figure 1, CAL-MAPF could have multiple unloading ports and we call each unloading port and cache around the port a group. Each group is independent in task queue, cache grids and agents. Agents in one group can only accept tasks from this group and use cache grids of this group. Algorithm 1 illustrates the operation of one group. Task assigner uses a state machine to control agent status and how they transfer to another based on different situations. When any agent reaches its target location, we must check all agents, update their status, and assign new target locations.
The state machine is illustrated in Figure 2, and the pseudocode of Algorithm 1 outlines how an agent transitions through five states while executing the low-level MAPF algorithm for one group. The task assigner monitors and updates the status of agents based on the item’s location, the state of the cache, and the agents’ current status. This algorithm ensures the efficient completion of all tasks by orchestrating the agents’ paths and their interactions with the cache and unloading port.
Figure 3: The average frequency of agent wait actions on each map grid with 256 agents. It is evident that CAL-MAPF experiences more severe traffic congestion compared to the baseline.
@@
Then, we evaluate whether other agents can alter their status. For agents in Status 0, heading to shelves, we examine if they have reached the shelves and gotten items. We then check if they can do a write operation in the cache and apply a cache replacement policy to select a cache grid. Thus, Status 0 may transition either to Status 2 (write to cache) or Status 3 (proceed directly to unloading port) (Algorithm 2 Lines 9-14). If they have not yet arrived, we check whether they can retrieve an item from the cache, which would change their status to 1 (read cache) (Algorithm 2 Lines 15-18). We do the same status changes for agents in Status 3 and Status 4 as we describe in  Figure 2 (Algorithm 2 Lines 19-28).
Figure 4: Makespan (Bar chart, lower is better) and Cache Hit Rate (Line chart, higher is better). LRU, FIFO and RANDOM represent CAL-MAPF with different cache replacement policies. NONE represents Lifelong MAPF without cache.
In Algorithm 3, we detail the process of reading or inserting (writing) a new item into cache grids using a cache replacement policy. For the read operation, it is necessary to verify if the item exists in the caches and whether there is an existing write lock on the cache grid (Algorithm 3 Lines 7-14). For the write operation, we first check whether the item is already in the cache area or if another agent plans to insert the same item into the caches (Algorithm 3 Lines 15-19). Then, we determine if an available cache exists for inserting the item (Algorithm 3 Lines 20-31).
@@ V Experimental Results
We use Lifelong MAPF without cache as a baseline, which aligns with our problem definition. To evaluate performance, we compare CAL-MAPF against Lifelong MAPF, utilizing LaCAM as the low-level MAPF solver for generating collision-free paths. Both CAL-MAPF and Lifelong MAPF were implemented in C++, building upon parts of the existing LaCAM codebase. [^1] The LaCAM source code is publicly accessible at https://github.com/Kei18/lacam. Our implementation is available at https://github.com/HarukiMoriarty/CAL-MAPF. All experiments were conducted on a system running Ubuntu 20.04.1, equipped with an AMD Ryzen 3990X 64-Core CPU and 64GB RAM at 2133 MHz.
@@ V-A Test Settings
As shown in Figure 1, we designed a warehouse map (27x71) with caches based on our problem definition. The map is adapted from the warehouse map of MAPF benchmark [2]. It includes 1600 shelf grids, a maximum of 80 cache grids and 4 unloading ports. We tested the CAL-MAPF in both multi-port and single-port scenarios as depicted in Figure 1. The multi-port scenario has 4 working groups of unloading ports and cache grids, each with a maximum of 20 cache grids. The single-port scenario maintains the same number of total cache grids and agents as the multi-port but only has one unloading port. Each shelf grid represents a unique kind of item, and we randomly assign an index to each shelf grid. The maximum cache-to-shelf ratio is 5%. We test all scenarios with different cache numbers {16, 32, 48, 64, 80} by deleting some cache grids in Figure 1. We have also chosen various total numbers of agents, $\{4,8,16,32,64,128,192,256\}$. In the multi-port scenario, each group is allocated an equal number of agents $\{1,2,4,8,16,32,48,64\}$.
Since we can expect the distribution of the task queue could significantly affect the performance of cache design, we designed three input task distributions to test CAL-MAPF, including: (1) $M$-$K$ distribution (MK): For any consecutive subarray of length $M$ in task queue $Q$, there are at most $K$ different kinds of items. This distribution is inspired by [37], where LRU has been proven to have a bound on cache miss rate and to be better than FIFO. This distribution can also represent that there are only several items people purchase daily, and some items may become very popular at one moment, replacing the previously popular items. (2) 7:2:1 distribution (Zhang): there are 70% kinds of items with only a 10% appearance probability in the task queue, 20% kinds of items with a 20% probability, and 10% with a 70% appearance probability [48]. (3) Real Data Distribution (RDD): we obtain public warehouse data from Kaggle222kaggle.com/datasets/felixzhao/productdemandforecasting, build a probability distribution based on item frequency of Kaggle data, and generate tasks from this probability distribution.
Figure 5: Makespan (Bar chart, lower is better) and Cache Hit Rate (Line chart, higher is better). As the number of caches increases, CAL-MAPF’s performance improves. In the MK distribution, CAL-MAPF outperforms the baseline in most test settings. For the Real and Zhang distributions, CAL-MAPF’s performance eventually surpasses the baseline. However, with a limited number of caches or with large number of agents, CAL-MAPF shows poor performance in both cache hit rate and makespan within the Real and Zhang distributions.
@@
We randomly generate $Q$ for all distributions with a total length of 1,000. For the MK distribution, we select $M=200$ and $K$ values of $\{1,20,40,80,120,160\}$. We allocate a total of 10 seconds for the low-level MAPF solver to find a collision-free solution for all 1,000 tasks in $Q$. We show performance using makespan which is the completion time when the last task has been accomplished. For CAL-MAPF, we use three different cache replacement policies: Least Recently Used (LRU), First-In-First-Out (FIFO), and RANDOM. We use NONE to represent the baseline method.

### v2

@@ Abstract
Multi-Agent Path Finding (MAPF), which involves finding collision-free paths for multiple robots, is crucial in various applications. Lifelong MAPF, where targets are reassigned to agents as soon as they complete their initial targets, offers a more accurate approximation of real-world warehouse planning. In this paper, we present a novel mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new type of map grid called cache for temporary item storage and replacement, and designed a locking mechanism for it to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. We identified three main factors significantly impacting CAL-MAPF performance through experimentation: suitable input task distribution, high cache hit rate, and smooth traffic. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions, maps, and agent configurations.
@@ I INTRODUCTION
Automated warehouses represent a multibillion-dollar industry led by corporations such as Amazon, Alibaba, Ocado, and inVia. These facilities deploy hundreds of robots to transport goods from one location to another [1]. Planning collision-free solutions for hundreds of robots is a key component of these automated warehouses, and can be simplified to a multi-agent path finding problem (MAPF) [2]. The MAPF problem requires planning collision-free paths for multiple agents from their start locations to pre-assigned target locations in a known environment while minimizing a given cost function. Many algorithms have been developed to solve this problem optimally and suboptimally, such as Conflict Based Search (CBS) [3], $M^{*}$[4] and Enhanced CBS (ECBS) [5].
@@
Although many studies focus on developing effective algorithms, there is also interest in improving the throughput of automated warehouses within the Lifelong MAPF framework by optimizing warehouse layouts. Prominent approaches include Fishbone design [12], DSAGE [13], and GGO [14]. However, these works mainly focus on static storage design and do not consider input distributions with specific patterns that may change over time. To address this problem, some studies have explored intelligent decision-making strategies [15, 16, 17]. Notable examples include the Kiva system and various policies, such as the velocity-based policy [16], and mixed-integer optimization and heuristic methods [18]. However, these methods are designed for specific warehouse systems, making it difficult to apply them to different warehouse types. They also require implementing complex rearrangement policies for the entire warehouse storage.
In this paper, we present a dynamic cache delivery mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), inspired by cache design, a foundational element extensively utilized in computer architecture, databases, and various other domains. This cache mechanism is adaptable to various Lifelong MAPF algorithms, warehouse storage strategies, and incoming task distributions. We have developed a novel map grid type that allows items to be temporarily stored and replaced. Additionally, we design a lock mechanism to enhance the stability of the planning solution. We have evaluated this cache mechanism using different cache replacement policies and a range of input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: input task distribution, cache hit rate, and map design. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.
@@ II Problem Definition
Lifelong Multi-Agent Path Finding (Lifelong MAPF) problem presents unique challenges due to the dynamic and unending nature of the tasks. Following the terminology in the MAPF community, we use agents to refer to robots. Let $I=\{1,2,\cdots,N\}$ denote a set of $N$ agents. $G=(V,E)$ represents an undirected graph, where each vertex $v\in V$ represents a possible location of an agent in the workspace, and each edge $e\in E$ is a unit-length edge between two vertices that moves an agent from one vertex to the other. Self-loop edges are allowed, which represent “wait-in-place” actions. Each agent $i\in I$ has a unique start location $s_{i}\in V$. Assume there is an task queue $Q=[q_{1},q_{2},...]$. For each task $q_{i}\in Q$, $q_{i}$ represents a certain item located in a reachable vertex in $V$ where at least one agent can find a path to the vertex. Each item must be delivered to a vertex in $V$ as the user specifies. These tasks are generated from a specific probability distribution to enhance realism.
@@
We study an on-line setting in which incoming tasks are not known in advance. It is assumed that there is a task assigner (external to low-level MAPF algorithm) from which agents can request target locations throughout the system’s operation. The task assigner determines the specific target locations for agents based on tasks in $Q$. In our paper, this task assigner is considered naive and will return the target location based on the first unfinished task in $Q$. Each agent $i\in I$ starts from a start location $s_{i}\in V$ and its target location depends on the task assigner. As agents complete their paths at their assigned goals, they are immediately assigned new targets, reflecting warehouse operations’ perpetual and iterative nature. Each action of agents, waiting in place or moving to an adjacent vertex, takes a unit of time. An agent’s path, $p^{i}=[v_{0}^{i},v_{1}^{i},\ldots,v_{T_{i}}^{i}]$, tracks the sequence of vertices traversed by agent $i$ from its start to its target.
As shown in  Figure 1, we primarily focus on 2D warehouse layout maps and categorize all non-obstacle map grids into three types: blue grids $S=\{s_{i}|s_{i}\in V\}$ for shelves that store items, white grids $W=\{w_{i}|w_{i}\in V\}$ for normal aisle grids, and green grids $U=\{u_{i}|u_{i}\in V\}$ for unloading ports where agents deliver items to complete a task. Agents can enter $S$ positions only from $W$ positions. Direct transitions between two positions in $S$ are not allowed.
@@
We will also involve a new type of map grid: caches. These additional vertices, designated as $C=\{c_{i}|c_{i}\in V\}$ in purple, serve as interim storage areas to streamline the completion of tasks by reducing the travel time of the agents retrieving items. To the best of our knowledge, no previous MAPF works have explored this new cache mechanism, so we adopt the following simplifying assumptions: (1) The warehouse is considered to have an unlimited supply of items, eliminating worries about stock running out. (2) Agents have infinite capacity to carry items. However, they are limited to transporting one type of item at a time, enabling them to move any quantity of that specific item without restriction. (3) Items removed or evicted from the cache are treated as if they vanish immediately.
@@ III-A MAPF
(One-Shot) MAPF, which has been proved an NP-hard problem with optimality [6], has a long history [19]. It has inspired a wide range of solutions for its related challenges. Decoupled strategies, as outlined in [19, 20, 21], approach the problem by independently planning paths for each agent before integrating these paths. In contrast, coupled approaches [22, 23] devise a unified plan for all agents simultaneously. There also exist dynamically coupled methods [3, 24] that consider agents that are planned independently at first and then together only when needed in order to resolve agent-agent conflicts. Among these, Conflict-Based Search (CBS) algorithm [3] stands out as a centralized, optimal solution for MAPF, with several bounded-suboptimal variants such as ECBS [5] and EECBS [25]. Some suboptimal MAPF solvers, such as Prioritized Planning (PP) [26, 19], PBS [27], PIBT [10] and their variant methods [28, 29, 11, 30] exhibit better scalability and efficiency.
@@ III-B Lifelong MAPF
Compared to the MAPF problem, Lifelong MAPF continuously assigns new target locations to agents once they have reached their current targets. It is not necessary for agents to arrive at their targets simultaneously in Lifelong MAPF. There are mainly three kinds of methods: solving the problem as a whole [31], using MAPF methods but replanning all paths at each specified length timestep [8, 10], and replanning only when agents reach their current target locations and are assigned new targets [32, 33, 11, 30]. Among them, MAPF-LNS [9, 29] and LaCAM [11, 30] are the leading methods with high performance in task execution.
@@ III-C Cache
The cache [34] serves as a crucial hardware component designed to temporarily store data, enhancing the efficiency of future request processing. Its main goal is to speed up access to frequently used data, thus minimizing dependence on slower storage disks. Popular caching policies include Least Recently Used (LRU) [35] and First-In-First-Out (FIFO) [36], with studies indicating LRU’s superiority over FIFO [37, 38, 39]. The caching concept is widely applied in various fields, such as Databases [40] and Content Delivery Networks (CDN) [41].
@@ III-D Warehouse Storage Strategy
Automated Storage and Retrieval Systems (AS/RS) have attracted attention for their potential to enhance warehouse efficiency and reduce operational costs [42, 43]. Various strategies for assigning items to storage locations have been widely adopted and evaluated [43, 44]. The random storage policy allocates each item type to a randomly chosen storage location and offers high space utilization [45]. The policy for storage at the closest open location places the arriving unit of items at the nearest available storage location, minimizing the immediate travel distance [46, 47]. The turnover-based storage policy assigns storage locations to each item based on their demand frequency [16, 17].
@@ IV-B Algorithm
As shown in Algorithms 1, 2 and 3, the whole procedure is mainly based on the task assigner, who encounters new tasks and decides what locations the agents need to visit. As illustrated in Figure 1, CAL-MAPF could have multiple unloading ports and we call each unloading port and cache around the port a group. Each group is independent in task queue, cache grids, and agents. Agents in one group can only accept tasks from this group and use cache grids of this group. Algorithm 1 illustrates the operation of a group. Task assigner uses a state machine to control agent status and how they transfer to another based on different situations. When any agent reaches its target location, we must check all agents, update their status, and assign new target locations.
The state machine is illustrated in Figure 2, and the pseudocode of Algorithm 1 outlines how an agent transitions through five states while performing the low-level MAPF algorithm for a group. The task assigner monitors and updates the status of agents based on the item’s location, the state of the cache, and the agents’ current status. This algorithm ensures the efficient completion of all tasks by orchestrating the agents’ paths and their interactions with the cache and unloading port.
Figure 3: The average frequency of agent wait actions on each map grid with 256 agents under Zhang distribution. It is evident that both CAL-MAPF and baseline experience severe traffic congestion as agent number is large enough.
@@
Then we evaluate whether other agents can alter their status. For agents in status 0, heading to shelves, we examine if they have reached the shelves and gotten items. We then check if they can do a write operation in the cache and apply a cache replacement policy to select a cache grid. Thus, status 0 may transition to status 2 (write to cache) or Status 3 (proceed directly to unloading port) (Algorithm 2 lines 9-14). If they have not yet arrived, we check whether they can retrieve an item from the cache, which would change their status to 1 (read cache) (Algorithm 2 Lines 15-18). We do the same status changes for agents in Status 3 and Status 4 as we describe in  Figure 2 (Algorithm 2 Lines 19-28).
Figure 4: Makespan (Bar chart, lower is better) and Cache Hit Rate (Line chart, higher is better). LRU, FIFO, and RANDOM represent CAL-MAPF with different cache replacement policies. NONE represents Lifelong MAPF without cache.
In Algorithm 3, we detail the process of reading or insertion (writing) a new item into cache grids using a cache replacement policy. For the read operation, it is necessary to verify if the item exists in the caches and if there is an existing write lock on the cache grid (Algorithm 3 lines 7-14). For the write operation, we first check whether the item is already in the cache area or if another agent plans to insert the same item into the caches (Algorithm 3 lines 15-19). Then, we determine if an available cache exists for inserting the item (Algorithm 3 Lines 20-31).
@@ V Experimental Results
We use Lifelong MAPF without cache as a baseline, which aligns with our problem definition. To evaluate performance, we compare CAL-MAPF with Lifelong MAPF, using LaCAM as the low-level MAPF solver for generating collision-free paths. Both CAL-MAPF and Lifelong MAPF were implemented in C++, building on parts of the existing LaCAM codebase [^1] The LaCAM source code is publicly accessible at https://github.com/Kei18/lacam. Our implementation is available at https://github.com/HarukiMoriarty/CAL-MAPF. All experiments were conducted on a system running Ubuntu 20.04.1, equipped with an AMD Ryzen 3990X 64-core CPU and 64GB RAM at 2133 MHz.
@@ V-A Test Settings
As shown in Figure 1, we designed a warehouse map (27x71) with caches based on our problem definition. The map is adapted from the warehouse map of MAPF benchmark [2]. It includes 1600 shelf grids, a maximum of 80 cache grids and 4 unloading ports. We tested CAL-MAPF in both multi-port and single-port scenarios, as depicted in Figure 1. The multi-port scenario has 4 working groups of unloading ports and cache grids, each with a maximum of 20 cache grids. The single-port scenario maintains the same number of total cache grids and agents as the multi-port but only has one unloading port. Each shelf grid represents a unique kind of item, and we randomly assign an index to each shelf grid. The maximum cache-to-shelf ratio is 5%. We test all scenarios with different cache numbers {16, 32, 48, 64, 80} by deleting some cache grids in Figure 1. We have also chosen various total numbers of agents, $\{4,8,16,32,64,128,192,256\}$. In the multi-port scenario, each group is assigned an equal number of agents $\{1,2,4,8,16,32,48,64\}$.
Since we can expect the distribution of the task queue could significantly affect the performance of cache design, we designed three input task distributions to test CAL-MAPF, including: (1) $M$-$K$ distribution (MK): For any consecutive subarray of length $M$ in the task queue $Q$, there are at most $K$ different kinds of item. This distribution is inspired by [37], where LRU has been proven to have a limit on the cache miss rate and to be better than FIFO. This distribution can also represent that there are only several items people purchase daily, and some items may become very popular at one time, replacing previously popular items. (2) 7:2:1 distribution (Zhang): there are 70% kinds of items with only a 10% appearance probability in the task queue, 20% kinds of items with a 20% probability, and 10% with a 70% appearance probability [48]. (3) Real Data Distribution (RDD): we obtain data from Kaggle’s public warehouse222kaggle.com/datasets/felixzhao/productdemandforecasting, build a probability distribution based on the frequency of Kaggle data items and generate tasks from this probability distribution.
Figure 5: Makespan (Bar chart, lower is better) and Cache Hit Rate (Line chart, higher is better). As the number of caches increases, CAL-MAPF’s cache hit rate and makespan performance improve. In the MK, Real, and Zhang distributions, CAL-MAPF surpasses the baseline in most test settings. However, it is observable that as the number of agents increases, the improvement offered by CAL-MAPF diminishes. Additionally, an abnormal increase in makespan is observed at baseline under the Zhang distribution with 256 agents.
@@
We randomly generate agent start locations and $Q$ for all distributions with a total length of 1,000. For the MK distribution, we select $M=200$ and $K$ values of $\{1,20,40,80,120,160\}$. We allocate a total of 10 seconds for the low-level MAPF solver to find a collision-free solution for all 1,000 tasks in $Q$. We show performance using makespan which is the completion time when the last task has been accomplished. For CAL-MAPF, we use three different cache replacement policies. Least Recently Used (LRU), First-In-First-Out (FIFO), and RANDOM. We use NONE to represent the baseline method.